# CATArena: Evaluation of LLM Agents Through Iterative Tournament Competitions

### v1

@@ E.1 Global Learning Trend
Figure 4: Trend of global performance score $G_{i}^{n}$ in Gomoku, Hold’em, Bridge and Chess.
As shown in Figure 4, we present the trends of global performance scores $G_{i}^{n}$ across four games, revealing distinct performance patterns for different models. In many cases, agents experience a sharp decline in performance during an intermediate round, which we interpret as a learning failure. Typically, such failures are recovered in the following round.
E.2 Detailed Learning Score
We list the detailed score of global learning, counter-adpatation learning and self-improvement in table 91011 respectively.
Table 9: Global Learning with Group-wise Average Rankings.
Models Avg. Ranking Gomoku Hold’em Bridge Chess
Standard Variant Standard Variant Standard Variant Standard Variant Standard Variant
Minimal
Claude-4-Sonnet 2.50 2.75 -0.447 -0.156 0.118 0.110 0.174 0.047 -0.170 0.018
Deepseek-Chat 2.75 2.75 0.027 0.077 0.010 -0.022 0.000 0.000 0.000 0.049
Doubao-Seed 4.75 4.50 -0.192 -0.302 -0.035 0.000 -0.033 -0.516 -0.337 0.034
Gemini-2.5-Pro 3.75 2.75 -0.066 0.173 0.020 0.078 -0.195 -0.049 -0.147 0.003
GPT-5 3.50 3.75 0.062 -0.019 0.102 -0.050 -0.095 0.293 -0.525 -0.000
Qwen3-Coder 3.75 4.50 -0.523 -0.089 0.038 0.003 0.032 -0.230 -0.187 -0.532
Commercial
best ADK 2.25 3.75 0.075 -0.022 0.073 0.030 0.295 0.361 -0.110 -0.342
Claude-Code 3.75 2.75 -0.322 0.194 0.100 0.105 -0.240 -0.139 -0.158 -0.226
CodeX 2.75 3.00 0.454 -0.095 0.050 0.067 0.098 0.285 0.033 0.064
Gemini-CLI 2.25 4.00 0.260 0.172 0.050 -0.007 0.254 -0.286 0.395 -0.154
Qwen-Coder 3.75 1.25 -0.054 0.536 0.058 0.105 0.157 0.556 -0.204 0.039
Table 10: Counter-adaptation Score with Group-wise Average Rankings.
Games Avg. Ranking Gomoku Hold’em Bridge Chess
Standard Variant Standard Variant Standard Variant Standard Variant Standard Variant
Minimal
Claude-4-Sonnet 3.75 3.75 -0.096 -0.075 0.001 -0.023 0.005 0.128 -0.042 -0.075
Deepseek-Chat 2.75 2.00 0.008 0.088 0.004 0.021 0.000 0.000 0.000 0.100
Doubao-Seed 4.25 4.50 0.063 -0.196 -0.023 -0.061 -0.008 -0.133 -0.196 0.083
Gemini-2.5-Pro 2.75 2.75 0.354 0.192 0.014 0.097 -0.132 -0.238 -0.038 0.021
GPT-5 4.75 3.25 0.038 -0.012 -0.086 0.019 -0.098 0.052 -0.154 -0.025
Qwen3-Coder 2.75 4.75 -0.092 0.029 0.000 -0.080 0.037 -0.008 0.025 -0.167
Commercial
best ADK 2.50 3.00 0.260 0.000 0.018 0.039 0.127 0.238 -0.094 -0.104
Claude-Code 3.25 1.88 -0.042 0.083 0.091 0.194 0.060 0.081 -0.104 0.047
CodeX 3.50 3.50 0.104 -0.089 -0.034 0.033 0.023 -0.090 0.010 0.052
Gemini-CLI 2.75 4.25 0.120 -0.021 0.031 -0.083 -0.027 -0.169 0.188 0.010
Qwen-Coder 3.00 2.38 -0.130 0.193 0.041 0.077 0.092 0.081 -0.062 -0.021
Table 11: Self-improvement Score with Group-wise Average Rankings.
Games Avg. Ranking Gomoku Hold’em Bridge Chess
Standard Variant Standard Variant Standard Variant Standard Variant Standard Variant
Minimal
Claude-4-Sonnet 2.75 2.50 -0.103 -0.894 0.949 0.517 0.858 0.766 -0.848 0.478
Deepseek-Chat 3.25 3.38 0.000 0.893 -0.949 -0.747 0.000 0.000 0.000 0.000
Doubao-Seed 3.75 4.38 0.141 -0.686 0.075 0.000 -0.202 -0.775 -0.894 0.000
Gemini-2.5-Pro 3.63 2.00 0.400 0.775 -0.758 0.894 -0.598 -0.240 -0.775 0.913
GPT-5 4.88 4.50 -0.897 -0.949 -0.205 -0.050 -0.767 0.485 -0.775 -0.390
Qwen3-Coder 2.75 4.25 -0.400 0.161 0.668 0.202 0.473 -0.400 -0.258 -0.775
Commercial
best ADK 2.25 3.50 0.400 0.956 0.835 -0.614 0.738 0.546 -0.207 -0.726
Claude-Code 4.00 2.25 -0.230 0.969 0.346 0.904 0.113 0.537 -0.730 -0.225
CodeX 2.00 3.25 0.763 -0.763 -0.090 0.602 0.784 0.316 0.424 0.316
Gemini-CLI 3.75 4.00 -0.183 -0.356 -0.176 0.000 0.000 -0.995 0.811 -0.193
Qwen-Coder 3.00 2.00 0.632 0.717 0.826 0.705 0.641 0.677 -0.944 0.000
The trend on four games are rather different. In general, the commercial model group consistently demonstrates superior global learning capability, where the advantage is particularly evident in complex strategy games like Chess or Gomoku variant.
Despite most minimal agents fail to learn well on complex games, we still find that Claude-4-Sonnet significantly surpass the rivals on standard games. However, the Claude-4-Sonnet still lacks behind on some cases like Gomoku, indicating that current LLMs agentic ablilities are still limited by the framework, where the commericial agents optimize workflow for their specific models to achieve the best results.
In simple games such as Hold’em, a larger proportion of agents exhibit positive learning scores, whereas in complex games like Chess, the prevalence of negative scores increases markedly. This trend suggests that current agents still face significant limitations in learning complex strategies.
E.3 Behavioral Changes Induced by Learning
For each game, we randomly select 80-100 intermediate states from the agents’ rival history and require the agent or LLM-player to choose the next move for each state. To ensure clarity in our writing, we uniformly refer to these intermediate states as endgame throughout the paper. Please note that endgame here is not limited to the final stages of the game; samples are taken from early, middle, and late stages as well.
We visualize the action consistency among the first two rounds in four games’s endgame in Figure 5.
From the matrix, we observe that, in general, agents tend to learn the strategies of other agents in the first round (lower left part vs. upper left part). Specifically, Doubao-Seed and DeepSeek-Chat simply copy Claude-4-Sonnet’s strategy in Holdem. Additionally, the learning trend is more pronounced in simpler games (Holdem vs. Chess).
For simpler games like Holdem, agent strategies in the second round are more similar to those in the first round (lower right part vs. upper left part), while for more difficult games like Chess, the trend is reversed. This observation is consistent with our findings on $\text{Trend}_{\text{mean}}$ in Table 5.
(a) Holdem.
(b) Gomoku.
(c) Bridge.
(d) Chess.
Figure 5: Action consistency between round 1 and round 2 agents’ code on endgames.
Appendix F Comparison between Agent and LLM-Player
Table 12: Comparison of match outcomes between each agent and its corresponding LLM-Player. Each value indicates the agent’s win rate when competing against the LLM that powers it.
Agent VS LLM Gomoku Hold’em Bridge Chess
Standard Variant Standard Variant Standard Variant Standard Variant
Claude-4-Sonnet 1.00 1.00 0.00 0.00 0.45 1.00 0.88 0.75
Deepseek-Chat 0.50 0.00 0.40 0.30 0.00 0.00 0.00 0.00
Doubao-Seed 0.50 0.25 0.00 0.00 0.00 0.10 0.00 0.00
Gemini-2.5-Pro 0.00 0.25 0.00 0.00 1.00 1.00 0.13 0.50
GPT-5 0.00 0.50 0.00 0.50 1.00 0.00 0.00 0.00
Qwen3-Coder 1.00 0.00 0.50 0.50 1.00 0.00 0.50 0.50
Claude-Code 1.00 1.00 0.00 0.00 0.70 0.85 0.75 0.63
CodeX 0.00 0.50 0.20 0.40 1.00 0.38 1.00 1.00
Gemini-CLI 0.00 0.00 0.10 0.40 0.00 1.00 0.38 0.50
Qwen-Coder 1.00 1.00 0.50 0.00 1.00 0.00 0.56 0.63
We compare the strategies of agents’ code and its corresponding LLM-Player on four games in table 12.
Interestingly, there is no strong correlation between the performance of the agents’ code and that of their underlying models.
In games with strong strategic elements, such as Gomoku and Chess, some agents’ code significantly outperforms their corresponding LLM-Player, indicating that the code implementation is able to better leverage game rules and strategies. For example, the agent developed by claude-4-sonnet achieves a 100% win rate against its LLM-Player in both standard and variant Gomoku, and also demonstrates a high win rate in Chess and Bridge. This suggests that the strategies implemented in the code are superior to the large model’s direct reasoning performance as a player in these games. In contrast, the agents developed by doubao-seed and deepseek-chat struggle to defeat their respective models. However, in Hold’em, agents generally have lower win rates than the LLM-Player, possibly because the LLM-Player performs better in games with more psychological tactics, which are difficult to simulate with code while can be summarized by context learning.
We further visualize the action consistency between agents’ code and LLM-palyers in Figure 6.
We find that, in most games, the actions of LLM-players tend to resemble each other, and the strategies implemented in agents’ code also exhibit high similarity among themselves. However, the code-based strategies and the plain reasoning of the same model often differ substantially. The only exception is Bridge, where we find numerous cases in which both LLM-players and code agents exhibit low consistency with human decisions. Considering that Bridge allows for a certain degree of decision freedom and its bidding rules are not strictly unified, we attribute this phenomenon to the intrinsic characteristics of the game. Similar observations are also reported in other studies (Kita et al., 2024).
These findings further demonstrate that the strategies generated by the agents and those employed by the LLM-players are based on different approaches. This difference merits additional investigation in subsequent studies.
Appendix G Case Study on Code
G.1 Strategy of Agents
In Gomoku, strategies display clear stratification. Gemini and DeepSeek rely on random or near-random moves, while Claude, Doubao, GPT-5, and Qwen3 employ similar pattern-based evaluation with candidate filtering and Minimax search. Differences mainly lie in threat recognition, search control, and opening play: Claude and Doubao handle openings and forced moves more effectively, GPT-5 is steadier under time limits, Qwen3 remains balanced, whereas Gemini and DeepSeek are notably weaker.
The code similarity among agents in Texas Hold’em is relatively high, only DeepSeek employs a fully random strategy, while other models calculate winning probabilities based on the hand. On one hand, this is because the available actions in Texas Hold’em are limited to fold, call/raise, and check. On the other hand, the strategies for Texas Hold’em are relatively straightforward to implement, as both reasoning and code are primarily based on hand strength. As a result, the code can closely simulate the reasoning process.
In the case of Chess, DeepSeek relies on an external library (Stockfish), but fails to configure it correctly, resulting in unsuccessful development. Even after multiple development iterations, DeepSeek continues to use this library without resolving the configuration issues. We also find that Claude, Doubao, Gemini, GPT, and Qwen3 utilize a similar combination of heuristic piece and board evaluation, Minimax search, and alpha-beta pruning, which leads to similar behavior. There are slight differences in how each model evaluates the value of Chess pieces and the actions in endgame scenarios. Notably, Claude incorporates an opening book, which distinguishes it from the others and leads to better performance.
For Bridge, bidding and play strategies also stratify clearly. Qwen3 and Gemini rely on minimal logic, following random choices or only basic rules on High Card Points (HCP). By contrast, GPT-5, Doubao, and Claude incorporate structured evaluation, moving from total point counting (GPT-5) to multi-layered systems with suit quality, competitive actions, and signaling (Doubao and Claude). Despite these differences, all models share reliance on HCP as a core metric. Overall, Claude achieves the most complete integration of evaluation and play, Doubao is comparably advanced, GPT-5 remains simpler, while Gemini and Qwen3 lag behind.
(a) Holdem.
(b) Gomoku.
(c) Bridge.
(d) Chess.
Figure 6: Action consistency between agents’ code and LLM-Players on endgames.
Appendix H Results of ML Track
@@
Minimal
Claude-4-Sonnet 0.787 0.360 0.600 0.700 2.25
@@
The detailed results of agents’ performance on machine learning devlopments is shown in table 13.
In ML track, agents autonomously generate data, design code, train models, and deliver ML-based strategies in a GPU-enabled environment. The results of ML ability is provided in Appendix H. Most agents only implement basic models with limited training, resulting in smaller performance gaps and different rankings compared to the strategy track.
Appendix I Results of Multi-lingual Track
We list the detailed results on different languages in  14.
Table 14: Scores of agents on games with different languages, with variance analysis.
[…]
Minimal
Claude-4-Sonnet 1.000 0.250 0.250 0.125 0.360 0.640 0.000 0.069 1.000 0.250 0.250 0.125 0.106
Commercial
Claude-Code 1.000 0.250 0.250 0.125 0.200 0.020 0.780 0.105 1.000 0.250 0.250 0.125 0.118
Most agents achieve their highest scores in Python, while several models exhibit significant performance fluctuations in JS and Go. Qwen3-Coder demonstrates the most consistent results across all languages, with the lowest average variance, indicating strong cross-language adaptability. In contrast, models such as GPT-5 and Doubao-Seed show considerable differences between languages, reflecting limited generalization ability. Commercial agents also exhibit score differences across different programming languages. Considering that board game strategies are inherently language-agnostic tasks, the performance gaps observed in the multi-language track of CATArena indicate that current agents are not yet able to effectively abstract strategies into unified algorithms and implement them consistently across languages. Such algorithmic abstraction should be a key direction for the future development of agents.
Appendix J Cost and Code Complexity of Participants
Table 15: Comparison of model cost on four games of 1st round.
@@ Appendix I Results of Multi-lingual Track
Average across games
Total input tokens 1208689.625 36707.75 87115.25 464059.75 370266.5 239983.25 637555.875 - 342248.875 496871.125
Session tokens 38442.75 10668.5 18443.0 41083.0 31757.25 19460.0 30771.125 - 30445.25 27170.375
Total output tokens 17093.5 1193.25 13365.0 40106.0 17752.75 5707.0 11396.0 - 8021.375 11981.75
Total time (s) 568.989 21.203 270.910 471.917 332.034 185.803 256.603 - 308.907 459.155
Tools used 77.5 10.0 15.5 51.25 33.5 26.5 26.75 20.25 14.375 30.875
Valid lines of code 642.75 22.5 354.5 199.0 362.25 305.0 332.25 327.5 253.0 357.25
Avg thinking time (s) 0.3489 0.0012 0.7503 0.0044 0.5867 0.0260 0.4975 0.1792 0.0324 0.0050
Table 16: Comparison of model cost on four games of 2nd round.
@@
Average across games
Total input tokens 1570364.75 72464.75 297566.75 842630.5 992826.5 2566477.75 1743925.25 - 536505.75 445568.75
Session tokens 93916.75 8146.5 43761.5 109111.0 86704.0 143595.25 50515.0 - 44458.25 43107.75
Total output tokens 12182.5 1087.38 13830.88 21083.38 11294.13 8876.13 15005.63 - 8060.25 11583.75
Total time (s) 622.63 334.81 346.27 865.02 415.34 1025.73 601.21 - 536.66 487.98
Tools used 41.5 26.75 24.5 36.25 42.75 57.75 42.25 17.75 18.0 29.0
Valid lines of code 549.25 22.5 418.25 269.75 408.25 362.0 548.25 334.75 298.25 534.5
Avg thinking time (s) 0.7000 0.0013 0.5367 0.0029 0.4231 0.0911 0.1925 0.0585 0.4669 0.0057
Table 17: Comparison of model cost on four variant games of 1st round.
@@
Average across games
Total input tokens 873125.5 32763.25 94909.5 897345.25 267624.75 340063.0 886002.0 - 369823.25 235447.25
Complete tokens 34555.25 9821.5 23031.75 31329.25 17193.5 12377.75 34906.0 - 31263.5 24084.0
Total output tokens 14267.25 1154.5 9934.5 27090.5 13740.0 8697.5 11363.25 - 6911.75 7056.75
Total time (s) 369.3838 18.7773 218.2230 839.8785 377.5413 1863.3930 274.1518 - 190.5898 227.2668
Tools used 53.5 10.25 15.25 48.5 35.25 97.5 31.5 15.0 15.5 23.25
Valid lines of code 668.5 31.5 410.0 280.75 302.75 237.5 325.0 341.0 188.0 322.5
Avg thinking time (s) 0.2022 0.0012 0.0103 1.7729 0.3052 0.0036 0.4970 0.0245 0.0063 0.2025
Table 18: Comparison of model cost on four variant games of 2nd round.

### v2

@@ Appendix H Pairwise Code Similarity Heatmaps
(e) OIBench.
(f) SWE-Perf.
Figure 8: Pairwise similarity matrices between Round 1 and 2 code submissions across six tasks. For competitive games (a-d), similarity is measured by Action Consistency in endgame positions. For programming tasks (e-f), similarity is measured by CodeBLEU.
Appendix I Case Study of Agent Learning
To demonstrate how agents actively improve their strategies by analyzing other agents’ code and tournament reports, we present a detailed case study focusing on Claude-Code_A in Texas Hold’em Poker tournaments. This case clearly shows that score changes across rounds reflect the agent’s understanding and adaptation to opponent strategies.
Development Planning and Evidence of Learning
Analysis of the Round 3 development log shows that the agent explicitly planned to study other agents’ strategies. The planning log has the following entries:
1. "Read game project documentation and requirements" - completed
2. "Analyze tournament rules and mode" - ongoing
3. "Study last round tournament reports and AI strategies" - pending
4. "Examine existing AI code from round 2" - pending
Tool Use Summary
Table 10 summarizes all the learning-related tool use operations performed by claude-code_A during the development of its Round 3 code. The tool use logs show that the agent first located and read the tournament report file, extracted win statistics using Grep (codex_A: 35 wins, gemini-cli_B: 30 wins, claude-code_A: 10 wins), then listed all agents, and finally read the code of the top performers.
Table 10: Summary of learning-related tool use operations by claude-code_A (Round 3 development).
Tool Operation Purpose
Glob Search **/tournament_report_tourney_*.json Locate tournament report files
Glob Search **/history_tourney_*.json Locate tournament history files
Read Read tournament_report_tourney_075455.json Analyze agent performance
Read Read first 100 lines of report Quickly understand report structure
Grep Search for "wins": field Extract win counts for each agent
Glob List round_2/**/* Discover all agents in previous round
Read Read own previous solutions Review existing implementation
Read Read codex_A code Peer-strategy extraction (Top performer, 35 wins)
Read Read gemini-cli_B code Cross-agent logic synthesis (second performer, 30 wins)
Strategy Adaptation in Code
Based on the discovery that top-performing agents (codex_A) utilized aggressive betting patterns, the agent transitioned from a static probability-based model to a dynamic opponent-aware strategy. Comparing the code between rounds, it is evident that the agent made targeted improvements based on its analysis:
  • Round 2: Basic Chen formula evaluation, simple tournament strategy, no opponent modeling.
  • Round 3: Added opponent analysis (analyze_opponents()), improved decision formula incorporating opponent aggression, refined stack and position factors.
Example of newly added code in Round 3:
1def analyze_opponents(self, game_state: Dict[str, Any]) $\rightarrow$ Dict[str, float]:
2 """Analyze opponent tendencies based on action history"""
3 action_history = game_state.get('action_history', [])
4 # ... analyze aggression for each opponent ...
5 return opponent_stats
6
7# Use opponent analysis in decision making
8opponent_stats = self.analyze_opponents(game_state)
9avg_opponent_aggression = sum(opponent_stats.values()) / max(len(opponent_stats), 1)
10decision_strength = (
11 hand_strength *
12 (1 + position_factor * 0.4) *
13 stack_factor *
14 bubble_factor *
15 (1 + (1 - avg_opponent_aggression) * 0.2)
16)
These code modifications are targeted responses to the strategies employed by top-performing agents, enabling more effective adaptation and strategic improvement.
Reflections in Development Summary
After completing development, the agent’s summary explicitly compares its performance and highlights improvements:
Agent Self-Reflection Log Performance Analysis from Round 2 Based on the tournament results: - codex_A: 35 wins (best performer) - gemini-cli_B: 30 wins - claude-code_A: 10 wins (my previous version) My new AI improves on the previous version with: - Better preflop ranges and position play - Enhanced tournament strategy - More sophisticated opponent modeling - Improved risk management
Similar learning behavior is observed for codex_A and other agents. This case study clearly demonstrates that agents actively analyze tournament reports and top-performing code via explicit tool use operations. The sequence and content of these operations, combined with targeted code improvements and reflective summaries, confirm that score changes across rounds are a reliable indicator of the agent’s growing understanding and adaptation to competitor strategies.
This micro-level behavioral loop consists of performance benchmarking, code synthesis, and self-reflection. It provides a mechanistic explanation for the macro-level performance leaps observed in our evolutionary trajectories, as agents successfully internalize superior logic from their competitive environment.
Appendix J Robustness Analysis on Variant Tasks
This section details the variant settings and analyzes the impact of rule changes on agent performance. The primary objective is to demonstrate the robustness of the CATArena framework: even when agents cannot rely on memorized opening books or standard algorithms, our metric for evolutionary capability $S_{\text{evo}}$ remains a reliable indicator of their learning capability.
Variant Configurations
We introduce specific rule perturbations designed to disrupt established priors while preserving the core reasoning logic:
  • Gomoku: Introduce forbidden points (e.g., double-three restrictions) to constrain the search space.
  • Texas Hold’em: Alter probability landscapes by removing specific cards and inverting hand rank hierarchies.
  • Chess: Adopt Chess960 (Fischer Random Chess) and modified movement rules to neutralize opening databases.
  • Bridge: Implement a "Card Exchange" mechanism between teammates to disrupt standard signaling conventions.
  • OIBench: Shift the optimization objective from time complexity to space complexity, requiring memory-efficient refactoring rather than speed optimization.
Table 11: Results of variant tasks. We report the Base Performance ($S_{\text{base}}$) and Evolutionary Capability ($S_{\text{evo}}$) of Minimal and Commercial Agents across five variants of standard tasks. Best results in each column are bolded.
Agents Competitive Objective
Gomoku Hold’em Bridge Chess OIBench
$S_{\text{base}}$ $S_{\text{evo}}$ $S_{\text{base}}$ $S_{\text{evo}}$ $S_{\text{base}}$ $S_{\text{evo}}$ $S_{\text{base}}$ $S_{\text{evo}}$ $S_{\text{base}}$ $S_{\text{evo}}$
Minimal Claude 0.78 -0.100 0.13 +0.017 1.00 +0.031 0.65 -0.014 0.23 +0.111
DeepSeek 0.38 +0.063 0.00 -0.008 0.10 0.000 0.10 +0.020 0.07 -0.021
Doubao 0.73 -0.201 0.00 0.000 0.45 -0.155 0.10 +0.018 0.22 +0.181
Gemini 0.00 +0.159 0.00 +0.043 0.60 -0.011 0.90 -0.016 0.56 -0.028
GPT 0.76 -0.008 0.87 +0.014 0.10 +0.014 0.45 +0.004 0.66 +0.003
Qwen 0.36 +0.041 0.00 -0.008 0.76 -0.021 0.80 -0.158 0.41 -0.011
Commercial Best ADK Agent 0.75 +0.012 0.46 +0.010 0.00 +0.109 1.00 -0.182 0.75 +0.036
Claude Code 0.66 +0.093 0.00 +0.038 0.93 +0.024 0.44 -0.058 0.74 +0.019
Codex 0.69 -0.087 0.17 +0.030 0.50 +0.013 0.34 +0.029 0.00 0.000
Gemini CLI 0.19 -0.058 0.37 0.000 0.83 -0.191 0.38 -0.012 0.66 +0.160
Qwen Code 0.22 +0.146 0.00 +0.049 0.25 +0.144 0.34 +0.011 0.70 -0.042
Results and Discussion
Table 11 compares the performance metrics between standard and variant settings. We observe two critical phenomena:
  1. 1.
    Widening Gap in Baseline Performance $S_{\text{base}}$: Unlike standard tasks where scores are often tightly clustered, variant tasks exhibit higher variance in zero-shot rankings. This is because rule perturbations cause less capable agents to fail in generating valid or robust strategies, thereby amplifying the performance distinction between agents that rely on memorization versus those with genuine reasoning capabilities.
  2. 2.
    Consistency in Evolutionary Capability: Crucially, the evolutionary metric retains its effectiveness in distinguishing agent capabilities. Despite the altered baselines, the relative evolutionary trends align with those in standard tasks (where most capable agents achieve $S_{\text{evo}}>0$). This demonstrates that CATArena effectively captures the gradient of learning ability: it can reliably differentiate between agents that can adapt to new rules and those that cannot, proving that the evaluation of evolutionary potential is robust to task variations.
Appendix K Results of Multilingual Track
Table 12: Scores of agents on games across programming languages, with variance analysis.
[…]
Minimal Claude-4-Sonnet 1.000 0.250 0.250 0.125 0.360 0.640 0.000 0.069 1.000 0.250 0.250 0.125 0.106
Commercial Claude-Code 1.000 0.250 0.250 0.125 0.200 0.020 0.780 0.105 1.000 0.250 0.250 0.125 0.118
Table 12 details the performance of agents across five programming languages: Python, C++, Java, JavaScript, and Go. The primary goal of this track is to assess whether agents can abstract strategic logic independently of implementation syntax.
  1. 1.
    Domain-Specific Language Bias: We observe a strong correlation between task type and language performance. In competitive tasks, agents achieve peak performance in Python, consistent with the prevalence of Python in AI and game scripting. Conversely, in OIBench, agents implementing algorithms in C++ significantly outperform those using Python. This suggests that agents rely heavily on the statistical distributions of their pre-training data (where competitive programming is dominated by C++) rather than possessing a generalized coding capability.
  2. 2.
    Generalization Gaps: Significant performance degradation is observed in languages that are less aligned with the specific domain (e.g., JavaScript/Go for games). Qwen3-Coder demonstrates the most consistent results across all languages with the lowest average variance, indicating strong cross-language adaptability. In contrast, models such as GPT-5 and Doubao-Seed show considerable fluctuations, reflecting limited generalization ability.
Considering that the core logic for both games and algorithms is inherently language-agnostic, the observed performance gaps indicate a critical failure in algorithmic abstraction. Current agents struggle to decouple strategic reasoning from syntactic implementation, failing to transfer a known solution (e.g., a C++ algorithm) to a different language context (e.g., Python) without performance loss.
Appendix L Results of ML Track
@@ Appendix I Case Study of Agent Learning
Minimal Claude-4-Sonnet 0.787 0.360 0.600 0.700 2.25
@@
In the ML track, agents are tasked with an end-to-end machine learning pipeline: autonomously generating data, designing architectures, training models in a GPU-enabled environment, and deploying the model for inference. The results are summarized in Table 13. Unlike the Strategy Track where logic is explicit, the ML Track requires agents to manage a complex, multi-stage workflow.
  1. 1.
    Limited Technical Depth. Despite the availability of GPU resources, we observe that agents generally fail to leverage advanced deep learning techniques. Most agents only managed to implement basic models (e.g., simple MLPs or decision trees) with limited training epochs. Additionally, a subset of agents fail to implement a valid training loop entirely, submitting models with randomized parameters instead of learned weights.
  2. 2.
    Commercial vs. ADK Agent. A distinct capability gap exists in end-to-end development stability. All commercial agents successfully completed the full lifecycle—from data generation to model deployment. In contrast, ADK agents struggled with pipeline integration, frequently failing to produce a runnable submission due to environment configuration errors or logical breaks in the workflow.
  3. 3.
    Performance vs. Rule-Based Track. Currently, the performance of agents in the ML track significantly lags behind the rule-based strategies observed in the standard Strategy Track. This underperformance is attributed to the agents’ inability to autonomously debug and optimize complex ML workflows. Consequently, due to the low baseline quality and high failure rate, we did not conduct multi-round evolutionary experiments for this track. However, as agent capabilities in coding and scientific modeling improve, the ML track will become a critical frontier for evaluating self-evolving intelligence in CATArena.
Appendix M Agent performance on Hard Games: Pommerman
(a) Start Status (step=0)
(b) End Status (step=800)
Figure 9: One Pommerman round demo. All agents failed to generate functional strategies, resulting in aimless movement or inactivity.
Beyond the fundamental board games discussed in previous sections, we further investigate more sophisticated competitive environments that necessitate the ability to perceive complex scenarios and synthesize highly strategic code. Specifically, we utilize the Pommerman environment (Resnick et al., 2018), which was originally designed for multi-agent AI research and featured in the NeurIPS 2018 Competition Track [^3] https://neurips.cc/Conferences/2018/CompetitionTrack. Inspired by the classic Nintendo game Bomberman, Pommerman is played on a randomly generated symmetric $11\times 11$ grid. The environment contains both rigid walls (indestructible and impassable) and wooden walls (destructible by bombs). Each agent can execute one of six actions per turn: Stop, Up, Left, Down, Right, and Bomb, with the ultimate objective of being the sole survivor.
Zero-Shot Performance Failure $S_{\text{base}}$
The environment’s infrastructure comprises approximately 4,500 lines of Python code, while a proficient human-authored solution typically requires around 1,000 lines of strategic logic to achieve plausible gameplay. We task the agents with developing autonomous AI strategies for Pommerman; however, almost all agent-generated strategies proved ineffective during live evaluation. Our observations indicate that these results typically fall into three failure paradigms:
  • Random Walk: The agent moves aimlessly across the field without ever deploying bombs.
  • Self-Destruction: The agent successfully deploys bombs but fails to execute evasion maneuvers, resulting in self-elimination.
  • Inactivity: The agent fails to produce valid operations within the turn limit, remaining stationary throughout the match.
Code Analysis and Implications
The generated code ranged from 100 to 400 lines, relying mostly on random actions or rudimentary Breadth-First Search (BFS). While Claude attempted to implement modular functions for tactical value estimation, its code was riddled with logical bugs that led to runtime errors or inactivity.
These results highlight a critical cold start problem in complex environments: current LLMs cannot bridge the gap between understanding game rules and implementing a minimum viable strategy (MVP).Without a functional baseline, the evolutionary feedback loop cannot be initiated, validating our decision to exclude such high-complexity tasks from the main evolutionary experiments.
Figure 9 illustrates a typical match where agents fail to engage meaningfully. The blue(Claude’s) and green(OpenAI’s) pommerman stay still, the red(Gemini’s) one find a more valuable place and then become inactive, the yellow(Qwen’s) pommerman walk in random.
Appendix N Cost and Code Complexity of Participants
We list the agents’ cost and code statistics in Table 14 for first round development of standard games,  15 for second round development of standard games, 16 for first round development of variant games and  17 for second round development of variant games.
We can see that game development token costs show minimal variation, while differences are significant due to model changes. Claude (both minimal and code-based agents) consumes significantly more input tokens than competitors, exceeding the average by over 2 times, while Gemini generates notably more output tokens compared to other models. GPT-5 offers the best trade-off between cost and performance. Among all agents, second-round game development requires more input tokens, while output token growth remains marginal. In addition, commercial agents consistently use fewer tokens than their minimal-agent counterparts.
In terms of code complexity, agents driven by Claude-4 model consistently surpass other agents in both the number of effective lines of code developed and the time spent considering development strategies. We observe that its development strategies are more sophisticated. Additionally, the complexity of its code increases with each iteration, which indirectly demonstrates the model’s exceptional learning capabilities.
Table 14: Model cost statistics on standard tasks in Round 1.
@@
OIBench
Total input tokens 1101801 127528 1056516 8320878 193403 472132 1591886 - - -
Session tokens 1117511 130209 1135372 8387950 216516 479643 1603116 - - -
Total output tokens 15710 2681 78856 67072 23113 7511 11230 - - -
Total time (s) 663.28 144.24 1982.12 3603.76 659.12 346.27 529.58 - - -
Tools used 39 19 22 95 17 25 40 - - -
Valid lines of code 312 149 242 375 167 273 381 - 548 135
Avg thinking time (s) - - - - - - - - - -
SWE-Perf
Total input tokens 12566767 1233619 2867580 1109026 5945407 12246895 - - - -
Session tokens 12622336 1240596 2977490 1155080 6049146 12308897 - - - -
Total output tokens 55569 6977 109910 46054 103739 62002 - - - -
Total time (s) 5028.39 2187.02 3284.32 2661.64 4082.8 4318.23 - - - -
Tools used 123 23 31 14 76 138 - - - -
Valid lines of code 257 0 31 100 125 133 - - - -
Avg thinking time (s) - - - - - - - - - -
Table 15: Model cost statistics on standard tasks in Round 2.
@@
OIBench
Total input tokens 689268 56536 734892 445268 331248 434818 1667173 - - -
Session tokens 697796 57894 817803 491666 372662 441152 1674074 - - -
Total output tokens 8528 1358 82911 46398 41414 6334 6901 - - -
Total time (s) 499.51 68.95 2024.68 1868.17 1204.64 249.76 471.25 - - -
Tools used 29 14 34 23 22 25 46 - - -
Valid lines of code 278 0 398 271 339 306 419 - 780 420
Avg thinking time (s) - - - - - - - - - -
SWE-Perf
Total input tokens 6720993 633967 1735119 435355 4530952 16977184 - - - -
Session tokens 6738614 635214 1803480 506936 4593913 17070695 - - - -
Total output tokens 17621 1247 68361 71581 62961 93511 - - - -
Total time (s) 3939.86 263.16 2344.74 2377.89 3652.3 2546.19 - - - -
Tools used 79 15 18 17 48 156 - - - -
Valid lines of code 34 0 13 39 31 62 - - - -
Avg thinking time (s) - - - - - - - - - -
Table 16: Model cost statistics on variant tasks in Round 1.
@@
OIBench
Total input tokens 1360571 33086 7669553 339500 325721 721710 2035052 27754 - -
Complete tokens 1380136 35679 7786790 406925 373777 732945 2052561 28256 - -
Total output tokens 19565 2593 117237 67425 48056 11235 17509 502 - -
Total time (s) 2433.550 80.090 3607.680 3600.980 1344.090 834.490 541.540 56.040 - -
Tools used 44 8 127 27 23 34 46 7 - -
Valid lines of code 224 82 421 248 276 271 248 0 405 371
Avg thinking time (s) - - - - - - - - - -
Table 17: Model cost statistics on variant tasks in Round 2.